# ReQISC: A Reconfigurable Quantum Computer Microarchitecture and Compiler Co-Design

### v1

@@ 5.2.2. Template-based synthesis
The ReQISC compiler employs a “pre-synthesis followed by assembly” strategy for compiling programs in their high-level IRs. In the pre-synthesis stage, all of the circuits are transformed into $\mathrm{CCX}$-based circuits via $\mathrm{MCX}$ decomposition (barenco1995elementary). The compiler extracts all distinct 3Q IRs from the benchmark suite, ignoring specific qubit indices, and determines their minimal-#2Q $\mathrm{SU}(4)$ synthesis templates through multiple rounds of approximate synthesis. In the assembly stage, the compiler unrolls 3Q IRs in an input $\mathrm{CCX}$-based circuit into their optimal synthesis schemes building on the pre-synthesized template library. The equivalent circuit classes (ECC) of each IR and its $\mathrm{SU}(4)$ synthesis templates are derived, according to the self-invertibility and control-bit permutability of the 3Q IR. That enables further $\mathrm{SU}(4)$ gate count reduction when selectively assembling near-neighbor IRs from their alternative ECC templates, as two $\mathrm{SU}(4)$s acting on the same pair of qubits can be fused into one $\mathrm{SU}(4)$. Figure 8 illustrates this synthesis procedure of a circuit snippet with consecutive Toffoli and Peres gates. Since the distinct 3Q IRs in real-world programs are finite, this template-based approach is computationally efficient and calibration-friendly.
@@ 5.3.1. Calibration overhead
Herein we discuss how the ReSIC compiler overcomes the calibration issue. First, the number of distinct $\mathrm{SU}(4)$s introduced by template-based synthesis is negligible, as the distinct 3Q IR patterns are finite in real-world applications (wille2008revlib). It is the hierarchical synthesis pass that incurs appreciable calibration overhead through exhaustive local optimization, in which approximate synthesis inevitably induces more diverse $\mathrm{SU}(4)$s’ parameters. Therefore, employing the hierarchical synthesis pass means sacrificing calibration efficiency for greater #2Q reduction.
Moreover, naïvely applying our scheme to variational quantum programs (e.g., QAOA (farhi2014quantum), UCCSD (barkoutsos2018quantum)) which necessitate multiple runs with variational gate parameters is unrealistic as it requires continual calibration of variational $\mathrm{SU}(4)$s in experiments. Instead, the workload of reconfiguring variational $\mathrm{SU}(4)$ gates can be shifted to reconfiguring 1Q gates, by decomposing $\mathrm{SU}(4)$s into a gate set with fixed 2Q gates (e.g., $\mathrm{SQiSW}$, $\mathrm{B}$) and parametrized 1Q gates. These 1Q parameters can be easily calibrated with constant experimental overhead by means of the PMW protocol (chen2023compiling; gong2023robust; wei2024native; chen2025efficient), which allows for the implementation of any continuous 1Q gate without explicit calibration by simply tuning the phase shift of microwave pulses. While this compromise may result in a slightly higher 2Q gate count, it represents a necessary trade-off for executing variational algorithms.
@@ 5.3.2. Qubit routing
Even with sophisticated mapping algorithms, significant #2Q overhead is induced during qubit routing by the insertion of $\mathrm{SWAP}$ gates (li2019tackling; liu2022not; zhang2021time). Leveraging the expressiveness of $\mathrm{SU}(4)$, we develop a dedicated mapping algorithm, mirroring-SABRE, by incorporating the $\mathrm{SU}(4)$-aware $\mathrm{SWAP}$ search strategy into the SABRE (li2019tackling).
@@
Mirroring-SABRE inherits the basic thought of SABRE (li2019tackling). SABRE attempts to map 2Q gates layer by layer via extracting the “front layer” $F$, peeling executable gates and searching $\mathrm{SWAP}$ gates to minimize the heuristic cost $H$ of the unresolved front layer. $H$ involving the current front layer and a $\mathrm{SWAP}$ candidate is designed to minimize the topological distance between upcoming qubits and promote parallelism. In our proposed mirroring-SABRE, we additionally define the “last mapped layer” $L$ as the set of 2Q gates that has no succeeding ones within the DAG constructed by already mapped 2Q gates, such that the $\mathrm{SWAP}$ search process prioritizes $\mathrm{SWAP}$ gates that $L$ can absorb, as illustrated in Figure 9. We also define an initial heuristic cost function $H_{0}(F,\mathrm{DAG},\pi,D)$ before each $\mathrm{SWAP}$ search epoch with the similar calculation method relative to the original cost function $H(F,\mathrm{DAG},\pi,D,\mathrm{SWAP})$ from SABRE. If there is a $\mathrm{SWAP}$ candidate that can be absorbed by $L$ and also leads to a lower $H$ than $H_{0}$, it will be selected as a $\mathrm{SWAP}$ mirroring effect that induces no #2Q overhead. Otherwise, the $\mathrm{SWAP}$ search proceeds similarly to SABRE.
@@ 5.4.2. Implementation
The ReQISC compiler is developed in Python to be a self-contained quantum programming SDK. It also operationalizes the ReQISC microarchitecture, enabling the generation of executable $\mathrm{SU}(4)$-based circuits with control parameters applicable to specific hardwares. In case of no available hardware, the output circuits are expressed in $\{\mathrm{Can}(x,y,z),\mathrm{U3}(\theta,\phi,\lambda)\}$. The approximate synthesis functionality leverages modules of BQSKit (bqskit). A key feature is its configurable approach to the calibration-performance trade-off. We provide two practical compilation schemes: “ReQISC-Full”, which employs all optimization passes including the hierarchical synthesis for aggressive gate count reduction, and “ReQISC-Red”, which bypasses this pass for minimal calibration overhead.
@@ 6. Evaluation
We evaluate ReQISC against SOTA counterparts on a comprehensive benchmark suite, demonstrating its superiority spanning multiple hardware types and showing clear advantages in its microarchitecture, as well as in end-to-end logical-level (topology-agnostic) and topology-aware compilation. All experiments ran on an Apple M3 Max laptop.
@@ 6.1.2. Baselines
“Qiskit”, “TKet”, and “BQSKit” are three primary baselines for comparison. In Qiskit, the standard O3 transpile pass is used. In Tket, the standard compilation procedure is constructed by sequentially employing PauliSimp and FullPeepholeOptimise passes, in which PauliSimp is the built-in Hamiltonian simulation program optimization engine (cowtan2019phase). To demonstrate that ReQISC’s benefits stem from its co-designed compilation strategy, we also compare against three variants: “Qiskit-$\mathrm{SU}(4)$” and “TKet-$\mathrm{SU}(4)$”, with a pass that partitions and fuses 2Q gates into $\mathrm{SU}(4)$s appended to the standard Qiskit/TKet compilation procedure, and “BQSKit-$\mathrm{SU}(4)$”, which employs its compile pass by designating a customized $\{\mathrm{Can},\mathrm{U3}\}$ gate set.
Table 1. Logical-level compilation comparisons in terms of the reduction rates of #2Q, Depth2Q, and pulse duration $T$. Metrics within the merged column “Benchmarks info.” indicate the circuit-level and pulse-level metrics for quantum programs in its original $\mathrm{CNOT}$-based representation. $T$ is in unit $g^{-1}$. Red./Full. indicates ReQISC-Red/Full compiler. For near-identity $\mathrm{SU}(4)$ gates (occurring in qft, pf, qaoa and uccsd), the gate mirroring technique introduced in Section 3 is employed. Throughout we assume the $\mathrm{XY}$-coupled Hamiltonian where the baseline $\mathrm{CNOT}$ scheme has gate duration $\tau=\pi/\sqrt{2}g$.
Benchmarks info. Avg. reduction of #2Q (%) Avg. reduction of Depth2Q (%) Avg. reduction of Duration (%)
Category (#) #Qubit #2Q Depth2Q Duration $T$ Qiskit TKet BQSKit Red. Full. Qiskit TKet BQSKit Red. Full. Qiskit TKet BQSKit Red. Full.
alu (12) 5-6 15-192 12-166 26.7-368.8 1.54 1.72 8.12 29.83 32.01 1.85 2.20 9.20 26.41 28.40 1.85 2.20 9.20 52.41 54.48
bit_adder (13) 4-15 14-5.1k 13-3.8k 28.9-8.5k 9.43 9.70 13.81 33.01 40.38 8.21 8.36 11.33 22.75 30.45 8.21 8.36 11.33 52.57 56.14
comp (19) 5-6 13-155 12-134 26.7-297.7 1.13 6.66 6.15 27.56 32.50 1.29 7.11 3.86 21.91 26.25 1.29 7.11 3.86 47.17 50.44
encoding (9) 3-15 10-3.3k 10-2.5k 22.2-5.6k 5.66 7.78 13.20 36.49 38.88 5.66 7.86 12.40 33.94 36.46 5.66 7.86 12.40 57.15 59.14
grover (1) 9-9 288-288 228-228 506.5-506.5 0.00 0.00 17.36 44.44 44.44 0.00 0.00 20.61 30.70 30.70 0.00 0.00 20.61 53.69 53.69
hwb (12) 4-170 26-4.4k 25-3.5k 55.5-7.8k 1.28 3.62 9.95 25.89 31.80 1.20 4.09 10.69 27.18 34.17 1.20 4.09 10.69 51.52 58.11
modulo (8) 5-7 9-62 8-56 17.8-124.4 9.64 9.84 14.92 34.18 38.90 10.67 10.67 15.51 32.49 38.80 10.67 10.67 15.51 58.94 62.74
mult (3) 12-48 99-1.6k 60-293 133.3-650.9 0.00 0.25 0.00 16.05 16.05 0.00 0.24 0.00 11.53 11.53 0.00 0.24 0.00 41.00 41.00
pf (9) 10-30 60-1.8k 60-1.8k 133.3-4.0k 0.96 27.41 2.20 87.59 87.59 0.96 71.85 3.46 98.63 98.63 0.96 71.85 3.46 98.70 98.70
qaoa (9) 8-24 192-13.2k 60-1.2k 133.3-2.7k 0.00 0.00 0.00 58.47 58.47 0.00 0.00 0.00 70.96 70.96 0.00 0.00 0.00 80.78 80.78
qft (3) 8-32 56-612 26-122 57.8-271.0 0.00 0.00 0.00 50.00 50.00 0.00 0.00 0.00 50.00 50.00 0.00 0.00 0.00 50.77 50.77
rip_add (4) 12-62 81-481 68-393 151.1-873.0 0.00 6.21 12.37 31.06 49.57 0.00 0.72 11.86 30.82 39.07 0.00 0.72 11.86 53.75 58.69
square (3) 10-13 745-3.8k 547-3.0k 1.2k-6.6k 0.02 0.20 9.49 31.42 38.16 0.00 0.26 5.59 19.75 24.25 0.00 0.26 5.59 45.98 50.95
sym (6) 7-14 107-29.3k 63-22.1k 140.0-49.1k 3.42 3.86 9.67 30.16 35.42 0.62 1.31 4.67 18.93 23.81 0.62 1.31 4.67 45.92 50.75
tof (4) 5-19 18-102 15-78 33.3-173.3 0.00 4.46 0.00 30.29 30.29 0.00 3.51 0.00 12.38 12.38 0.00 3.51 0.00 39.39 39.39
uccsd (14) 8-14 630-14.4k 615-13.6k 1.4k-30.2k 23.46 64.03 6.71 76.50 84.11 23.49 67.53 7.48 78.45 85.63 23.49 67.53 7.48 84.02 89.25
urf (3) 8-9 9.0k-23.6k 7.3k-18.6k 16.2k-41.3k 0.91 4.47 6.81 20.61 25.90 0.83 4.66 3.30 14.34 20.77 0.83 4.66 3.30 42.71 48.55
@@ 6.1.3. Benchmarks
The benchmark suite comprises $132$ programs from $17$ diverse categories (alu, encoding, uccsd, etc.). Program sizes as indicated by #2Q in $\mathrm{CNOT}$-based circuit representation range from several to $10^{4}$. Most programs are selected from published benchmark suites (wille2008revlib; sivarajah2020t).
Notations of most benchmarks align with those used in published benchmark suites for which we refer readers to RevLib’s website and TKet’s benchmarking repo. For benchmarks whose high-level algorithmic description is in form of a .real file involving $\mathrm{MCX}$ gates, we convert them into $\mathrm{CCX}$-based circuits as the input for the ReQISC compiler and further decompose them into $\mathrm{CNOT}$-based circuits as the input for the baselines. For variational quantum programs, Pauli strings with coefficients are compiled into $\mathrm{SU}(4)$ gate sequences by the high-level ISA-independent Phoenix compiler (yang2025phoenix) before being used as input for the ReQISC compiler, while $\mathrm{CNOT}$-based circuits are fed to baselines.
Table 2. Synthesis cost comparison in terms of gate duration $\tau$ ($g^{-1}$) achieved by the ReQISC microarchitecture. The Haar-random average cost (number of basis gates for synthesizing arbitrary $\mathrm{SU}(4)$ gates) for $\mathrm{CNOT}$, $\mathrm{iSWAP}$, $\mathrm{SQiSW}$, $\mathrm{B}$ gates are 3, 3, 2.21, and 2, respectively.
Coupling type $\mathrm{XY}$ coupling $\mathrm{XX}$ coupling Random

### v2

@@ 5.2.2. Template-based synthesis
The ReQISC compiler employs a “pre-synthesis followed by assembly” strategy for compiling programs in their high-level IRs. In the pre-synthesis stage, all of the circuits are transformed into $\mathrm{CCX}$-based circuits via $\mathrm{MCX}$ decomposition (Barenco et al., 1995). The compiler extracts all distinct 3Q IRs from the benchmark suite, ignoring specific qubit indices, and determines their minimal-#2Q $\mathrm{SU}(4)$ synthesis templates through multiple rounds of approximate synthesis. In the assembly stage, the compiler unrolls 3Q IRs in an input $\mathrm{CCX}$-based circuit into their optimal synthesis schemes building on the pre-synthesized template library. The equivalent circuit classes (ECC) of each IR and its $\mathrm{SU}(4)$ synthesis templates are derived, according to the self-invertibility and control-bit permutability of the 3Q IR. This enables further $\mathrm{SU}(4)$ gate count reduction when selectively assembling near-neighbor IRs from their alternative ECC templates, as two $\mathrm{SU}(4)$s acting on the same pair of qubits can be fused into one $\mathrm{SU}(4)$. Figure 10 illustrates this synthesis procedure of a circuit snippet with consecutive Toffoli and Peres gates. Since the distinct 3Q IRs in real-world programs are finite, this template-based approach is computationally efficient and calibration-friendly.
@@ 5.3.1. Calibration overhead
Herein we discuss how the ReQISC compiler overcomes the calibration issue. First, the number of distinct $\mathrm{SU}(4)$s introduced by template-based synthesis is negligible, as the distinct 3Q IR patterns are finite in real-world applications (Wille et al., 2008). It is the hierarchical synthesis pass that incurs appreciable calibration overhead through exhaustive local optimization, in which approximate synthesis inevitably induces more diverse $\mathrm{SU}(4)$s’ parameters. Therefore, employing the hierarchical synthesis pass means sacrificing calibration efficiency for greater #2Q reduction.
Moreover, naïvely applying our scheme to variational quantum programs (e.g., QAOA (Farhi et al., 2014), UCCSD (Barkoutsos et al., 2018)) which necessitate multiple runs with variational gate parameters is unrealistic as it requires continual calibration of variational $\mathrm{SU}(4)$s in experiments. Instead, the workload of reconfiguring variational $\mathrm{SU}(4)$ gates can be shifted to reconfiguring 1Q gates, by decomposing $\mathrm{SU}(4)$s into a gate set with fixed 2Q gates (e.g., $\mathrm{SQiSW}$, $\mathrm{B}$) and parametrized 1Q gates. These 1Q parameters can be easily calibrated with constant experimental overhead by means of the PMW protocol (Chen et al., 2023; Gong and Choi, 2023; Wei et al., 2024; Chen et al., 2025), which allows for the implementation of any continuous 1Q gate without explicit calibration by simply tuning the phase shift of microwave pulses. While this compromise may result in a slightly higher 2Q gate count, it represents a necessary trade-off for executing variational algorithms.
@@ 5.3.2. Qubit routing
Even with sophisticated mapping algorithms, significant #2Q overhead is induced during qubit routing by the insertion of $\mathrm{SWAP}$ gates (Li et al., 2019; Liu et al., 2022; Zhang et al., 2021). Leveraging the expressiveness of $\mathrm{SU}(4)$, we develop a dedicated mapping algorithm, mirroring-SABRE, by incorporating the $\mathrm{SU}(4)$-aware $\mathrm{SWAP}$ search strategy into the SABRE (Li et al., 2019).
@@
Mirroring-SABRE inherits the basic thought of SABRE (Li et al., 2019). SABRE attempts to map 2Q gates layer by layer via extracting the “front layer” $F$, peeling executable gates and searching $\mathrm{SWAP}$ gates to minimize the heuristic cost $H$ of the unresolved front layer. $H$ involving the current front layer and a $\mathrm{SWAP}$ candidate is designed to minimize the topological distance between upcoming qubits and promote parallelism. In our proposed mirroring-SABRE, we additionally define the “last mapped layer” $L$ as the set of 2Q gates that has no succeeding ones within the DAG constructed by already mapped 2Q gates, such that the $\mathrm{SWAP}$ search process prioritizes $\mathrm{SWAP}$ gates that $L$ can absorb, as illustrated in Figure 11. We also define an initial heuristic cost function $H_{0}(F,\mathrm{DAG},\pi,D)$ before each $\mathrm{SWAP}$ search epoch with a similar calculation method
$H_{0}:=\frac{1}{\left|{F}\right|}\sum\nolimits_{g\in F}D_{\pi[g.q_{1}],\pi[g.q_{2}]}+\frac{W}{\left|{E}\right|}\sum\nolimits_{g\in E}D_{\pi[g.{q_{1}}],\pi[g.{q_{2}}]},$
relative to the original cost function $H(F,\mathrm{DAG},\pi,D,\mathrm{SWAP})$ from SABRE. The $\mathrm{SWAP}$ mirroring mechanism prioritizes $\mathrm{SWAP}$ candidates that can be absorbed by $L$ while simultaneously reducing the heuristic cost ($H<H_{0}$), thus inducing no #2Q overhead. If no such candidate meets these criteria, the $\mathrm{SWAP}$ search proceeds according to the $\mathrm{SABRE}$ heuristic.
@@ 5.4.2. Implementation
The ReQISC compiler is developed in Python to be a self-contained quantum programming SDK. It also operationalizes the ReQISC microarchitecture, enabling the generation of executable $\mathrm{SU}(4)$-based circuits with control parameters applicable to specific hardwares. In case of no available hardware, the output circuits are expressed in $\{\mathrm{Can}(x,y,z),\mathrm{U3}(\theta,\phi,\lambda)\}$. The approximate synthesis functionality leverages modules of BQSKit (Younis et al., 2021). A key feature is its configurable approach to the calibration-performance trade-off. We provide two practical compilation schemes: “ReQISC-Full”, which employs all optimization passes including the hierarchical synthesis for aggressive gate count reduction, and “ReQISC-Eff”, which bypasses this pass for minimal calibration overhead.
Table 1. Benchmark suite characteristics. Metrics per category include qubit count range (#Qubit), 2Q gate count (#2Q), 2Q depth (Depth2Q), and original circuit duration ($T$). Duration is reported in units of $g^{-1}$, assuming a baseline $\mathrm{CNOT}$ gate duration of $\pi/\sqrt{2}g$.
Category (#) #Qubit #2Q Depth2Q Duration $T$
alu (12) 5-6 15-192 12-166 26.7-368.8
bit_adder (13) 4-15 14-5.1k 13-3.8k 28.9-8.5k
comparator (19) 5-6 13-155 12-134 26.7-297.7
encoding (9) 3-15 10-3.3k 10-2.5k 22.2-5.6k
grover (1) 9-9 288-288 228-228 506.5-506.5
hwb (12) 4-170 26-4.4k 25-3.5k 55.5-7.8k
modulo (8) 5-7 9-62 8-56 17.8-124.4
mult (3) 12-48 99-1.6k 60-293 133.3-650.9
pf (9) 10-30 60-1.8k 60-1.8k 133.3-4.0k
qaoa (9) 8-24 192-13.2k 60-1.2k 133.3-2.7k
qft (3) 8-32 56-612 26-122 57.8-271
ripple_add (4) 12-62 81-481 68-393 151.1-873
square (3) 10-13 745-3.8k 547-3.0k 1.2k-6.6k
sym (6) 7-14 107-29.3k 63-22.1k 140-49.1k
tof (4) 5-19 18-102 15-78 33.3-173.3
uccsd (14) 8-14 630-14.4k 615-13.6k 1.4k-30.2k
urf (3) 8-9 9.0k-23.6k 7.3k-18.6k 16.2k-41.3k
Overall (132) 3-170 9-3.3k 8-2.2k 17.8-49.1k
@@ 6. Evaluation
Table 2. Logical-level compilation comparisons in terms of the reduction rates of #2Q, Depth2Q, and pulse duration. Eff./Full. denote ReQISC-Eff/Full. For near-identity $\mathrm{SU}(4)$ gates (in qft, pf, qaoa and uccsd), the gate mirroring technique introduced in Section 4.3 is applied. Durations are evaluated under the XYXY-coupled Hamiltonian, with baseline $\mathrm{CNOT}$ duration $\tau=\pi/\sqrt{2}g$.
Benchmarks Average reduction of #2Q (%) Average reduction of Depth2Q (%) Average reduction of Duration (%)
Category (#) Qiskit TKet BQSKit Eff. Full. Qiskit TKet BQSKit Eff. Full. Qiskit TKet BQSKit Eff. Full.
alu (12) 1.54 1.72 8.12 29.83 32.01 1.85 2.20 9.20 26.41 28.40 1.85 2.20 9.20 52.41 54.48
bit_adder (13) 9.43 9.70 13.81 33.01 40.38 8.21 8.36 11.33 22.75 30.45 8.21 8.36 11.33 52.57 56.14
comparator (19) 1.13 6.66 6.15 27.56 32.50 1.29 7.11 3.86 21.91 26.25 1.29 7.11 3.86 47.17 50.44
encoding (9) 5.66 7.78 13.20 36.49 38.88 5.66 7.86 12.40 33.94 36.46 5.66 7.86 12.40 57.15 59.14
grover (1) 0.00 0.00 17.36 44.44 44.44 0.00 0.00 20.61 30.70 30.70 0.00 0.00 20.61 53.69 53.69
hwb (12) 1.28 3.62 9.95 25.89 31.80 1.20 4.09 10.69 27.18 34.17 1.20 4.09 10.69 51.52 58.11
modulo (8) 9.64 9.84 14.92 34.18 38.90 10.67 10.67 15.51 32.49 38.80 10.67 10.67 15.51 58.94 62.74
mult (3) 0.00 0.25 0.00 16.05 16.05 0.00 0.24 0.00 11.53 11.53 0.00 0.24 0.00 41.00 41.00
pf (9) 0.96 27.41 2.20 87.59 87.59 0.96 71.85 3.46 98.63 98.63 0.96 71.85 3.46 98.70 98.70
qaoa (9) 0.00 0.00 0.00 58.47 58.47 0.00 0.00 0.00 70.96 70.96 0.00 0.00 0.00 80.78 80.78
qft (3) 0.00 0.00 0.00 50.00 50.00 0.00 0.00 0.00 50.00 50.00 0.00 0.00 0.00 50.77 50.77
ripple_add (4) 0.00 6.21 12.37 31.06 49.57 0.00 0.72 11.86 30.82 39.07 0.00 0.72 11.86 53.75 58.69
square (3) 0.02 0.20 9.49 31.42 38.16 0.00 0.26 5.59 19.75 24.25 0.00 0.26 5.59 45.98 50.95
sym (6) 3.42 3.86 9.67 30.16 35.42 0.62 1.31 4.67 18.93 23.81 0.62 1.31 4.67 45.92 50.75
tof (4) 0.00 4.46 0.00 30.29 30.29 0.00 3.51 0.00 12.38 12.38 0.00 3.51 0.00 39.39 39.39
uccsd (14) 23.46 64.03 6.71 76.50 84.11 23.49 67.53 7.48 78.45 85.63 23.49 67.53 7.48 84.02 89.25
urf (3) 0.91 4.47 6.81 20.61 25.90 0.83 4.66 3.30 14.34 20.77 0.83 4.66 3.30 42.71 48.55
We evaluate ReQISC against state-of-the-art (SOTA) counterparts on a comprehensive benchmark suite, demonstrating its superiority across multiple hardware types and showing clear advantages in its microarchitecture, as well as in end-to-end logical-level (topology-agnostic) and topology-aware compilation. All experiments were run on an Apple M3 Max laptop.
@@ 6.1.2. Baselines
“Qiskit”, “TKet”, and “BQSKit” are three primary baselines for comparison. In Qiskit, the standard O3 transpile pass is used. In Tket, the standard compilation procedure is constructed by sequentially employing PauliSimp and FullPeepholeOptimise passes, in which PauliSimp is the built-in Hamiltonian simulation program optimization engine (Cowtan et al., 2019). To demonstrate that ReQISC’s benefits stem from its co-designed compilation strategy, we also compare against three variants: “Qiskit-$\mathrm{SU}(4)$” and “TKet-$\mathrm{SU}(4)$”, with a pass that partitions and fuses 2Q gates into $\mathrm{SU}(4)$s appended to the standard Qiskit/TKet compilation procedure, and “BQSKit-$\mathrm{SU}(4)$”, which employs its compile pass by designating a customized $\{\mathrm{Can},\mathrm{U3}\}$ gate set.
@@ 6.1.3. Benchmarks
The benchmark suite comprises $132$ programs from $17$ categories (alu, encoding, etc.) shown in Table 1. Program sizes indicated by #2Q in $\mathrm{CNOT}$-based circuit representation range from tens to $10^{4}$. Most programs are selected from published benchmark suites (Wille et al., 2008; Sivarajah et al., 2020).
Notations of most benchmarks align with those used in published benchmark suites for which we refer readers to RevLib and TKet’s benchmarking repo. For benchmarks whose high-level algorithmic description is provided in .real format, featuring $\mathrm{MCX}$ gates as basic subroutines, we convert them into $\mathrm{CCX}$-based circuits as the input for the ReQISC compiler and further decompose them into $\mathrm{CNOT}$-based circuits as the input for baselines. For variational quantum programs, Pauli strings with coefficients are compiled into $\mathrm{SU}(4)$ gate sequences by the high-level ISA-independent PHOENIX compiler (Yang et al., 2025) before being used as the input for ReQISC, while $\mathrm{CNOT}$-based circuits are fed to baselines.
Table 3. Synthesis cost comparison in terms of gate duration $\tau$ ($g^{-1}$) achieved by ReQISC. Haar-random average costs (number of 2Q basis gates required for arbitrary $\mathrm{SU}(4)$ synthesis) for $\mathrm{CNOT}$, $\mathrm{iSWAP}$, $\mathrm{SQiSW}$, $\mathrm{B}$ gates are 3, 3, 2.21, and 2, respectively.
Coupling $\mathrm{XY}$ coupling $\mathrm{XX}$ coupling Random